OpenStudy (anonymous):

Jorge's rectangular garden was 17 feet long and had a perimeter of 54 feet. He made his garden 5 feet longer. What is the area of his new garden? A.64 square feet B.220 square feet C.85 square feet D.22 square feet

OpenStudy (aric200):

The answer is B . 220 square feet 17*2=34 54-34=20/2=10 feet wide 17+5=22 ft long* 10 wide= 220

OpenStudy (aric200):

@tatyy